bigrob012:

how do you change -x-2y=-5 to ymx+b

Shadow:

\[y = mx + b\] This is the form we want to change into. \[-x -2y = -5\] As you can see in the above formula, x is on the right side of the equals sign, so we should move that over. Whatever we do to one side, we must do to the other, to keep the equation 'true.' Think of it like saying, both Johnny and Sarah have the same amount of pie, that is what the equals sign represents, both sides. If I do something to Johnny's slice of pie, in order to keep the equation true, I must also do the same thing to Sarah's slice of pie. So if I take one spoonful, I will take the same spoonful from the other side. So I need to move -x, so I'll do the opposite action of its value. Since it's negative, I'll add x to both sides \[-x -2y + x = -5 + x\] \[-x + x = 0\] That cancels out, so we get \[-2y = x -5\] Now we have to isolate y, and since \[-2y = -2 \times y\], the opposite action would be to divide by -2, removing both the 2 multiplier from y, isolating it, and changing it to be positive, which it is in the format of y = mx + b \[\frac{ -2y = x -5 }{ -2 } \rightarrow y = - \frac{ 1 }{ 2 }x + \frac{ 5 }{ 2 }\]

Shadow:

Let me know if you have any questions. Depending on your teacher, you may need to convert 5/2 fraction into a decimal.

bigrob012:

thank you for the help

bigrob012:

okay thanks

Shadow:

You're welcome